A balloon Rises from rest with a constant acceleration g/8. A stone is released from it when it has rising to height h. The time taken by the stone to reach the ground is....???

Grade:Select Grade

1 Answers

Rohit kumar
6 years ago
Velocity acquired by the balloon at height h ,V=√2ah =√2*g/8*h =√gh/2Taking downward motion of a stone released from balloon at height H up to surface of ground. we have,u =-√gh/2;a=g;s=h,t=tAs , s =ut+1/2at^2=h =-2√gh/2*t+1/2gt2Or t^2 -√h/g*t-2h/g=0On solving,t=2√h/g or -√h/gAs negative time is not accounted , hence t =2√h/gAnswer